OpenStudy (anonymous):

(2a^-2b)^-3 over 5a^2b^4

OpenStudy (anonymous):

Is this what you're working with? \[(2a^{-2}b)^{-3} / 5a^2b^4\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So remember your rules for exponents.. \[(y^bx^a)^c = x^{ac}y^{bc}\] \[x^a * x^b = x^{a+b}\] \[x^a/x^b = x^{a-b}\] See how far those get ya.